OpenStudy (anonymous):

please please help!! Find the slope of the line containing the given pair of points. If the line is undefined, state so. Please be sure to show all of your work. (-5, -2) and (-7, -2)

OpenStudy (saifoo.khan):

Slope is 0

OpenStudy (anonymous):

y = -2

OpenStudy (anonymous):

how did you get the answer y=-2 please :)

OpenStudy (anonymous):

or is it slope 0?

OpenStudy (saifoo.khan):

Coz the line crosses the y axis at 2. while the slope is 0.

OpenStudy (saifoo.khan):

y=x(0)+b y = b where b is the y intercept.

OpenStudy (anonymous):

Basically we have a line that has to go through two points that have y values of -2 so y has to always be -2 hence y = -2

OpenStudy (anonymous):

Props to you, nurse, for wanting to know how the answer was achieved. The general way is to first compute the slope. Slope is (y2 - y1) / (x2 - x1) -- rise over run, as they say. Try substituting the actual points into that equation and see what you get.

OpenStudy (anonymous):

thanks ktklown so (y2-y1)/(x2-x1)=1-1=0?

OpenStudy (anonymous):

x1, y1, and x2, y2 are the two points you were given.

OpenStudy (anonymous):

(-5, -2) and (-7, -2)

OpenStudy (anonymous):

ktklown I still get 0 is this correct? thanks so much for explaining and helping me:)

OpenStudy (anonymous):

Yep, it's 0, I just wanted to make sure you were using the right method :-)

OpenStudy (anonymous):

you are awesome!! thanks again for explaining!!!!

OpenStudy (anonymous):

Now the general form of what's called the "point slope" form of a line is this: (y - y1) = m (x - x1). In this case, x1 and y1 are any point on the line, and m is the slope you calculated in the first step.

OpenStudy (anonymous):

So you substitute in the x1, y1, and m, and you're left with an equation that just has a X and a Y -- and that's your line. You can use either of the points you were given as x1 and y1, just as long as you're consistent about which one you use.

OpenStudy (anonymous):

I have a few more questions if you have time to help??

OpenStudy (anonymous):

I have time for one more.

OpenStudy (anonymous):

Find the x- and y-intercepts for the equation. Please show all of your work. Then graph the equation and submit your graph -x + 2y = 4

OpenStudy (anonymous):

OK, this is actually easier. The x intercept is where the line intercepts the X-axis, and that happens when Y is equal to 0. Similarly, the Y-intercept is where X=0. So to find the X=intercept, set y=0, and solve. Give it a shot! (Show your work so I can look for mistakes.)

OpenStudy (anonymous):

-x=4 x=-4 (-4,0)??????

OpenStudy (anonymous):

YES! Nicely done!

OpenStudy (anonymous):

Now do the same to find the other intercept by setting x=0 instead.

OpenStudy (anonymous):

Im so excited!!!!! thank you!!!!!

OpenStudy (anonymous):

im stuck sorry

OpenStudy (anonymous):

Great, I'm excited for you too! Remember that, in theory, you could set either x or y to *any* value you wanted, and solve the equation. Every solution will give you *some* point on the line. You're just being asked to find two *special* points on the line, that is, the places where it touches the X or Y axis. But you could plug literally *any* number into the equation -- x=1934, or x=3423424, and it would give you a point on the line.

OpenStudy (anonymous):

That's ok. So, just rewrite the equation with x set to 0, and show me what you get.

OpenStudy (anonymous):

-x=0 x=-0 (-0,4)??

OpenStudy (anonymous):

So the equation is -x + 2y = 4 So if you substitute x=0 then you get -(0) + 2y = 4 Does that make sense?

OpenStudy (anonymous):

yes thanks so much

OpenStudy (anonymous):

so what do you get as y?

OpenStudy (anonymous):

y=0

OpenStudy (anonymous):

well I get 0 + 2y = 4 which is the same as 2y = 4 And if you divide both sides by 2 you get y = 2

OpenStudy (anonymous):

thank you!!! you've been a tremendous help!!!!